# VLA-Pruner: Temporal-Aware Dual-Level Visual Token Pruning for Efficient Vision-Language-Action Inference

### v2

@@ 4.1 Problem Formulation
We define the token pruning problem as follows: given a set of visual tokens with and a target subset size , the goal is to select a subset that preserves information critical for VLA inference. Formally, we define a mapping that selects an index subset with , and constructs the retained visual tokens as . The objective is to identify a mapping function that minimizes the difference in the model’s output before and after pruning:
where and . denotes the conditional generation probability of VLA model. denotes a loss function that measures the difference in the model’s output with and without pruning, and indicates the number of retained tokens. We conceptually decompose into vision-language prefill and action decode . Let denote all hidden representations of textual and visual tokens after prefilling. Informally,
where also drops the corresponding indices from the hidden visual representations . As noted in Motivation 2, optimizing with a single-level objective, i.e., semantic-level or action-level , leads to suboptimal pruning. To address this, we propose VLA-Pruner.
@@ 4.2 The Proposed VLA-Pruner
VLA-Pruner adopts a (a) dual-level token importance criterion that incorporates both semantic-level and action-level importance for visual token retention, and proposes a corresponding (b) dual-level token selection strategy to adaptively process dual-level relevance and redundancy for token retention. The overall pipeline is illustrated in Fig. 3.
(a) Dual-level Token Importance Criterion
For VLMs, previous works [8, 60] have demonstrated that prefill or text-to-vision attention scores can effectively quantify significance of vision tokens. In VLA case, as discussed before (Sec. 3.2), both semantic-level and action-level visual requirements should be satisfied. To this end, VLA-Pruner adopts a dual-level token importance criterion: (1) prefill attention scores (Eq. 2) to quantify semantic-level relevance and (2) action decode attention scores (Eq. 3) to quantify action-level importance. As discussed in Motivation 3, we build on the temporal continuity of VLA manipulation to estimate current action attention scores over image patches. This formulates a time-series forecasting task [5], where is estimated from historical trend . Classical training-free forecasting methods include averaging, autoregressive models, and exponential smoothing [6, 5, 52]. Among them, the Exponential Moving Average (EMA) [21, 15, 14] provides a simple yet effective way to estimate temporal trends. By emphasizing recent observations while discounting older ones, EMA reduces noise, captures temporal trends, and has been widely applied in domains such as finance, economics, meteorology, and robotic manipulation [17, 43, 10, 39]. Directly applying EMA to our task takes this form:
where is the smoothing factor, denotes the smoothed prediction at the previous timestep, and represents the corresponding observation result.
Decaying Window Average Mechanism. For VLA manipulation, where short-term context is more critical [55], we instead propose a simpler and more intuitive decaying window average mechanism. This method considers recent timesteps and applies an explicit exponential decay. Formally, we smooth action attention scores from previous steps, , and compute the estimated score for current step as their weighted sum:
@@
Temporal smoothing methods can reliably estimate action attention dynamic trends but may fail under abrupt changes, such as target switching. For instance, in the LIBERO-Long task “put both the alphabet soup and the tomato sauce in the basket”, action attention from putting the soup does not ground the tomato sauce well. Potential remedies include lightweight predictors for attention shifts or mechanisms to bypass pruning when large prediction–observation discrepancies occur. To ensure generalizability and flexibility, we instead tackle this fundamental challenge through our dual-level token selection strategy.
(b) Dual-Level Token Selection Strategy.
Based on the dual-level token importance criterion, an intuitive approach to retain tokens is: normalizing semantic and action attention scores ( and ), taking a weighted sum, and selecting the top- tokens by the weighted score. However, this scheme suffers from fundamental limitations. First, it introduces a sensitive weighting hyperparameter that must be carefully tuned, compromising robustness and generalizability. Second, it ignores redundancy both within and across stages. On one hand, as noted in prior work [2], attention-based pruning often retains similar tokens. On the other hand, score-level fusion tends to retain tokens that appear salient in both prefilling and decoding, yet not essential to either stage. For example, in a robotic arm, intermediate arm segments may appear salient across both stages, whereas the end-effector (critical for execution) can be undervalued. Finally, it implicitly assumes that the action attention estimate is always reliable, which may not hold under abrupt attention shifts (see Remark 1).
Patch-wise Combine-then-Filter. To avoid these limitations, we draw on the minimal-Redundancy-Maximal-Relevance (mRMR) principle [42], an information-theoretic paradigm for feature selection. mRMR proposes to combine both Max-Relevance and Min-Redundancy principles, reducing the redundant features while keeping the relevant features for the task. Inspired by this, our dual-level token selection strategy follows a patch-wise combine-then-filter mechanism. We organize the procedure into three phases:
@@
(2) Relevance-Maximization Pooling. We then combine these two subsets to form a candidate pool:
This subset constructs a candidate pool that contains tokens essential for dual-level information requirements, maximizing two-stage task relevance. Due to attention divergence (Motivation 2), is generally larger than .
(3) Redundancy-Minimization Filtering. This phase aims to reduce feature redundancy within the candidates to satisfy the target budget . We achieve this by maximizing the diversity of the final selected subset. Following prior work [2], we formulate it as a Max–Min Diversity Problem (MMDP) [45] within candidate pool , where the goal is to find a subset of size that maximizes the minimum pairwise distance between its elements:
@@
A solution for the MMDP problem 8 is a subset of that minimizes redundancy between elements. MMDP can be solved by several exact and heuristic methods [46, 38]. To facilitate efficiency, we use a greedy algorithm that iteratively add tokens whose minimum distance to the retained set is maximal, until reaching elements. To escape local optima, we initialize the first token with maximal second-nearest distance. We detail the step (3) and our overall token selection method in Algorithm 1 and 2 (see Appendix A).
@@
Step (3) shares similarities with DivPrune [2], which reduces redundancy through diversity-based pruning. However, DivPrune does not incorporate relevance maximization step, thus fail to identify pivotal tokens. By combining both Max-relevance and Min-redundancy, we retain token subsets that are both diverse and better matched to dual-level information requirements within VLA inference.
(c) Implementation Details.
VLA-Pruner prunes visual tokens at Transformer layer K within VLA model (K is set to 3 [8]). We select tokens using last-layer prefill attention and temporal smoothing of action-to-vision attention averaged over the latter half layers (to reduce noise) from past timesteps. Pruned tokens are dropped from the rest prefill layers. To adapt to flash-attention [12], adaptation techniques in SparseVLM [60] can be adopted. For decaying window average, window size and decay rate are empirically set as 3 and 0.8. We warm-start VLA-Pruner with steps to record sufficient action attention history. VLA-Pruner is a plug-and-play acceleration module for VLA architectures with action-to-vision cross-attention, a mechanism used by state-of-the-art VLA models [22, 24, 4], including (i) autoregressive policies [22, 7, 63, 63] (e.g., OpenVLA [22]), (ii) action-chunk decoders (OpenVLA-OFT [24]), and (iii) diffusion-head policies such as [4], for which we average action attention scores over its flow-matching steps. We provide analysis of computational complexity in Appendix A.2.
@@
5 Experiments
We evaluate VLA-Pruner in both simulation and real-world settings. In simulation setting, we evaluate VLA-Pruner on open-source VLA models: OpenVLA [22], OpenVLA-OFT [24] and [4], using the LIBERO benchmark [32] and SIMPLER environment [27]. All experiments are conducted on an NVIDIA RTX 4090 GPU.
@@ 5.1 Experimental Setup
Baselines. We compare VLA-Pruner with a comprehensive suite of training-free accelerators under given compute budgets. Baselines include state-of-the-art token pruning methods (FastV [8], SparseVLM [60], and DivPrune [2]) and a VLA-specific method VLA-Cache [55]. We provide more details of baselines and VLA models in Appendix B.1. EfficientVLA [56] adopts token prune strategies similar to prior work (using attention and diversity), we include results of it and more token pruning baselines in Appendix LABEL:app:more_baselines. Evaluation Protocol. All methods are evaluated under identical given pruning budgets. To ensure a comprehensive comparison, our experiments include diverse pruning ratios (50%, 75%, 87.5%). For VLA-Cache, we set the same token reuse ratio. Our evaluation metrics mainly include: task success rate (%), inference latency (ms) and FLOPs(T).
5.2 Evaluation Benchmarks
In this section, we introduce the evaluation benchmarks, with more details presented in Appendix B.2.
LIBERO.
The LIBERO benchmark [32] comprises four suites (Spatial, Object, Goal, and Long) that probe complementary aspects of manipulation generalization. Each suite contains 10 distinct tasks with 50 evaluation episodes per task (500 total episodes per suite). Following OpenVLA [22] and OpenVLA-OFT [24], we adopt their standard evaluation setup (action chunks as 8 for OpenVLA-OFT) and official weights to ensure comparability. To verify cross-architecture generality, we also include results of diffusion-head based model [4].(see Appendix LABEL:app:pi0_libero)
SIMPLER.
The SIMPLER simulators [27] provide two complementary settings aimed at closing sim-to-real gaps: visual matching (VM) and variant aggregation (VA). We report evaluation results on three tasks and their corresponding settings: Move Near, Pick Coke Can (under VA), and Open/Close Drawer (under VM). We choose these task–setting pairs where the vanilla OpenVLA attains non-trivial success rates, ensuring a reasonable comparison.
Real Robot Evaluation.
To assess VLA-Pruner’s performance in real-world scenarios, we conduct experiments using a 6-DoF xArm6 robotic arm equipped with a parallel gripper. We run four manipulation tasks involving diverse objects (100 trials for each task): (1) Can Stack: grasp the can and stack it on top of the box; (2) Cup Pour: pick up the cup and pour the small balls inside the cup into the box; (3) Cube Place: grasp the cube and place it into the box; (4) Bottle Place: grasp the bottle and place it into the box.
@@ 5.3 Main Results
We evaluate VLA-Pruner on OpenVLA (fine-tuned) and OpenVLA-OFT for LIBERO benchmark. Table 4.2 reports success rate on the four suites under token-retention ratios of 50%, 25%, and 12.5%. We also report a relative accuracy metric that measures pruning-induced error, defined as the performance preservation ratio (vanilla model as 100%). VLA-Pruner achieves the best success rate and relative accuracy across all settings among training-free accelerators. Even at a 12.5% retention ratio, it maintains 88.9% and 88.27% relative accuracy on OpenVLA and OpenVLA-OFT, respectively, outperforming baselines by up to 34.39%. At a 50% retention ratio, VLA-Pruner can even improve success rate, thanks to precise noise filtering that stabilizes policy execution. This is evident on LIBERO-Long, where it achieves 4.2% success rate improvement. DivPrune benefits VLM inference but degrades VLA performance since robot manipulation needs more localized visual details. VLA-Cache achieves competitive results due to temporal awareness but fails at high compression ratios because its text-to-vision attention selection strategy is coarse-grained. Fig. 4 visualizes more details of VLA-Pruner performance varying under different pruning ratios. We present results on in Appendix LABEL:app:pi0_libero.
@@
We evaluate real-world performance on a physical 6-DoF xArm6 arm controlled by OpenVLA-OFT. Fig. 5 reports results under 75% token-pruning ratio. VLA-Pruner attains the highest relative accuracy, underscoring its practicality for on-robot deployment.
@@
We report inference efficiency metrics in Table 4.2, including inference latency ( or ) and FLOPs(T). At same pruning ratios, VLA-Pruner achieves comparable speedups to visual token–pruning methods. Its latency is 5 ms higher than FastV (due to distance computations in redundancy filtering), yet it achieves significantly better performance. At the 25% retention ratio, VLA-Pruner outperforms token pruning baselines at 50% retention, while being 16% faster and using 20% fewer FLOPs. VLA-Pruner is faster and more compute-efficient than VLA-Cache. VLA-Pruner requires storing historical attention, but it introduces a negligible increase in Max GPU memory (see Appendix LABEL:app:more_results:memory).
@@ 5.4 More Results
To validate our (a) dual-level importance criterion and (b) patch-wise dual-level token selection strategy, we ablate three variants: (1) prefill-only, which selects tokens by prefill attention; (2) action-only, which selects tokens by action attention; and (3) score-fusion, which ranks tokens by the weighted sum of prefill and action attention scores, with a carefully fine-tuned weight. As shown in Fig. 6, prefill-only preserves high-level semantics but hinders fine-grained low-level control, whereas action-only improves short-horizon control but sacrifices semantic grounding and task planning, and thus neither maintains VLA performance well. Score-fusion remains inferior to the prefill-only setting. We also ablate the temporal decay window size in Fig. 6, which demonstrates the robustness of our method and reveals the short-term temporal continuity of VLA manipulation. When , it reduces to using the last timestep’s action attention to guide pruning, leading to performance degradation. We further ablate varying decay rate and pruning layer in the Appendix LABEL:app:more_results:ablations.
@@
We provide a qualitative case study of VLA-Pruner in Fig. 7. VLA-Pruner preserves essential localized visual information for precise action execution, while maintaining broad semantic understanding.

### v3

@@ 4.1 Problem Formulation
We define the VLA token pruning problem as follows: given a set of visual tokens with and a target subset size , the goal is to select a subset that preserves information critical for VLA inference. We define a mapping function selects an index subset with , and constructs the retained visual tokens as . The objective is to identify a that minimizes the difference in the model’s output before and after pruning:
where and . denotes the conditional generation probability of VLA model. measures the difference in the model’s output with and without pruning, and indicates the number of retained tokens. To formalize the intuition behind Motivation 2, we decompose into vision-language prefill and action decode based on hidden states :
Here, sparsifies both the input and corresponding hidden states . For VLA inference, relying on single-level objectives—either semantic-level or action-level —fails to preserve the joint probability . To address this, we introduce VLA-Pruner.
@@ 4.2 The Proposed VLA-Pruner
VLA-Pruner adopts a dual-objective importance criterion and a corresponding dual-level token selection strategy for VLA token pruning. The overall pipeline is illustrated in Figure 3 and detailed in Algorithm 1.
4.2.1 Dual-objective Importance Criterion
For VLMs, previous works (Chen et al., 2024; Zhang et al., 2024) have demonstrated that prefill or text-to-vision attention scores effectively quantify the importance of vision tokens. For VLA models, as discussed before (Sec. 3), both semantic-level and action-level visual requirements should be satisfied. Thus, VLA-Pruner adopts a dual-objective importance criterion: (i) prefill attention scores (Eq. 2) to quantify semantic-level relevance and (ii) action attention scores (Eq. 3) to quantify action-level importance.
Attention Estimation via Temporal Smoothing. As discussed in Motivation 3, we build on the temporal continuity of VLA manipulation to estimate current action attention scores over image patches. This formulates a time-series forecasting task (Box et al., 2015), where is estimated from historical trend . We adopt the Exponential Moving Average (EMA) technique (Hyndman and Athanasopoulos, 2018; Gardner Jr, 2006; Gardner, 1985) to ensure stable estimation and adaptability to dynamic shifts. EMA reduces noise, captures temporal trends by emphasizing recent observations while discounting older ones, and has been widely applied in domains such as economics, meteorology, and robotic manipulation (Contreras et al., 2003; Mhammed et al., 2016). Directly applying EMA to our task takes this form:
where is the smoothing factor, denotes the smoothed prediction at the previous timestep, and represents the corresponding observation result. For VLA manipulation, where short-term context is more critical (Xu et al., 2025), we adapt the standard EMA into a Decaying Window Average mechanism. Formally, we smooth action attention scores from previous steps and compute the estimated score for current step as their weighted average:
@@ 4.2.1 Dual-objective Importance Criterion
Temporal smoothing can reliably estimate action attention dynamic trends but may fail under abrupt changes, such as target switching. For instance, in the LIBERO-Long task “put both the alphabet soup and the tomato sauce in the basket”, attention from putting the soup does not ground the tomato sauce well. Potential remedies include lightweight predictors or mechanisms to bypass pruning when large discrepancies occur. To ensure generalizability and flexibility, we instead tackle this fundamental challenge through our dual-level token selection strategy.
4.2.2 Dual-Level Token Selection Strategy
Based on the dual-objective token importance criterion, an intuitive approach to retain tokens is: normalizing semantic and action attention scores ( and ), taking a weighted sum, and selecting the top- tokens by the weighted score. However, this scheme suffers from critical limitations. First, it introduces a sensitive weighting hyperparameter that must be carefully tuned, compromising robustness and generalizability. Second, it ignores redundancy both within and across stages. On one hand, as noted in prior work (Alvar et al., 2025), attention-based pruning often retains similar tokens. On the other hand, score-level fusion tends to retain tokens that appear salient in both prefilling and decoding, yet not essential to either stage. For example, in a robotic arm, intermediate arm segments may appear salient across both stages, whereas the end-effector (critical for execution) can be undervalued. Third, it implicitly assumes that the action attention estimate is always reliable, which may not hold under abrupt attention shifts (see Remark 1).
To avoid these limitations, we draw on the minimal-Redundancy-Maximal-Relevance (mRMR) principle (Peng et al., 2005), an information-theoretic paradigm for feature selection. mRMR proposes to combine both Max-Relevance and Min-Redundancy principles, keeping the relevant features for the task while reducing the redundant features. Inspired by this principle, our proposed dual-level token selection strategy follows a patch-wise Combine-then-Filter paradigm. We organize the procedure into three phases:
@@ 4.2.2 Dual-Level Token Selection Strategy
(2) Relevance-Maximization Pooling.
We then take the union of these two subsets to form a candidate pool:
This subset constructs a candidate pool that contains tokens essential for dual-level information requirements, maximizing two-stage task relevance.
Due to attention divergence (Motivation 2), is generally larger than .
(3) Redundancy-Minimization Filtering.
This phase aims to reduce feature redundancy within the candidates to satisfy the target budget . We achieve this by maximizing the diversity of the final selected subset. Following prior work (Alvar et al., 2025), we formulate it as a Max–Min Diversity Problem (MMDP) (Porumbel et al., 2011) within candidate pool , where the goal is to find a subset of size that maximizes the minimum pairwise distance between its elements:
@@
A solution for the MMDP problem 8 is a subset of that minimizes redundancy between elements. MMDP can be solved by several exact and heuristic methods (Resende et al., 2010; Martí et al., 2022). To facilitate efficiency, we use a greedy algorithm that iteratively add tokens whose minimum distance to the retained set is maximal, until reaching elements. To escape local optima, we initialize the first token with maximal second-nearest distance. We detail the step (3) in Algorithm 2 (see Appendix A).
@@
Step (3) shares similarities with DivPrune (Alvar et al., 2025), which reduces feature redundancy through diversity-based pruning. However, DivPrune does not incorporate the Max-relevance step, and thus fails to identify pivotal visual tokens. By combining both Max-relevance and Min-redundancy principles, we retain token subsets that are both diverse and better matched to dual-level visual information requirements for VLA inference.
4.3 Implementation Details.
VLA-Pruner prunes visual tokens at Transformer layer K (K set to 3 (Chen et al., 2024)). It selects tokens using last-layer prefill attention and temporal smoothing of action-to-vision attention averaged over the latter half layers. Pruned tokens are dropped from the rest prefill layers. For decaying window average, window size and decay rate are empirically set as 3 and 0.8. We warm-start VLA-Pruner with steps to record sufficient attention history. VLA-Pruner is training-free, serving as a plug-and-play acceleration module for VLA architectures with action-to-vision cross-attention, a mechanism used by state-of-the-art VLA models (Kim et al., 2024, 2025b; Black et al., 2024), including (i) autoregressive policies (Kim et al., 2024; Brohan et al., 2022; Zitkovich et al., 2023) (e.g., OpenVLA (Kim et al., 2024)), (ii) action-chunk decoders (OpenVLA-OFT (Kim et al., 2025b)), and (iii) diffusion-head policies such as (Black et al., 2024), for which we average action attention over flow-matching steps. We provide theoretical analysis of computational complexity in Appendix A.2.
5 Experiments
@@
We evaluate VLA-Pruner in both simulation and real-world settings. We evaluate VLA-Pruner on open-source VLA models: OpenVLA (Kim et al., 2024), OpenVLA-OFT (Kim et al., 2025b) and (Black et al., 2024). All experiments are conducted on an NVIDIA RTX 4090 GPU.
@@ 5.1 Experimental Setup
(1) Baselines.
We compare VLA-Pruner with a comprehensive suite of training-free accelerators under given compute budgets. Baselines include state-of-the-art token pruning methods (FastV (Chen et al., 2024), SparseVLM (Zhang et al., 2024), and DivPrune (Alvar et al., 2025)) and a VLA-specific method VLA-Cache (Xu et al., 2025). We provide more details of baselines and VLA models in Appendix B.1. EfficientVLA (Yang et al., 2025) uses token pruning strategies similar to prior work (using attention and diversity). We include results of it and HoloV (Zou et al., ) in Appendix LABEL:app:more_baselines.
(2) Evaluation Protocol.
All methods are evaluated under identical given pruning budgets. To ensure a comprehensive comparison, our experiments include diverse pruning ratios (50%, 75%, 87.5%). For VLA-Cache, we set the same token reuse ratio. Our evaluation metrics mainly include: task success rate (%), inference latency (ms) and FLOPs(T).
(3) Evaluation Benchmarks.
We introduce the evaluation benchmarks, with more details presented in Appendix B.2: LIBERO. The LIBERO benchmark (Liu et al., 2023a) comprises four suites (Spatial, Object, Goal, and Long) that probe complementary aspects of manipulation generalization. Each suite contains 10 distinct tasks with 50 evaluation episodes per task (500 total episodes per suite). Following OpenVLA (Kim et al., 2024) and OpenVLA-OFT (Kim et al., 2025b), we adopt their standard evaluation setup (action chunks as 8 for OpenVLA-OFT) and official weights to ensure comparability. To verify cross-architecture generality, we also include results of diffusion-head based model (Black et al., 2024).(see Appendix LABEL:app:pi0_libero). SIMPLER. The SIMPLER simulators (Li et al., 2024b) provide two complementary settings aimed at closing sim-to-real gaps: visual matching (VM) and variant aggregation (VA). We report evaluation results on three tasks and their corresponding settings: Move Near, Pick Coke Can (under VA), and Open/Close Drawer (under VM). We choose these task–setting pairs where the vanilla OpenVLA attains non-trivial success rates, ensuring a reasonable comparison. Real Robot Evaluation. To assess VLA-Pruner’s performance in real-world scenarios, we conduct experiments using a 6-DoF xArm6 robotic arm equipped with a parallel gripper. We run four manipulation tasks involving diverse objects (100 trials for each task): (1) Can Stack: grasp the can and stack it on top of the box; (2) Cup Pour: pick up the cup and pour the small balls inside the cup into the box; (3) Cube Place: grasp the cube and place it into the box; (4) Bottle Place: grasp the bottle and place it into the box.
[…]
@@ 5.2 Main Results
We evaluate VLA-Pruner on OpenVLA (fine-tuned) and OpenVLA-OFT for LIBERO benchmark. Table 5 reports success rate on the four suites under token-retention ratios of 50%, 25%, and 12.5%. We also report a relative accuracy metric that measures pruning-induced error, defined as the performance preservation ratio (vanilla model as 100%). VLA-Pruner achieves the best success rate and relative accuracy across all settings among training-free accelerators. Even at a 12.5% retention ratio, it maintains 88.9% and 88.27% relative accuracy on OpenVLA and OpenVLA-OFT, respectively, outperforming baselines by up to 34.39%. At a 50% retention ratio, VLA-Pruner can even improve success rate, thanks to precise noise filtering that stabilizes policy execution. This is evident on LIBERO-Long, where it achieves 4.2% success rate improvement. DivPrune benefits VLM inference but degrades VLA performance since robot manipulation needs more localized visual details. VLA-Cache achieves competitive results due to temporal awareness but fails at high compression ratios because its text-to-vision attention selection strategy is coarse-grained. Figure 4 visualizes more details of VLA-Pruner performance varying under different pruning ratios. We present results on in Appendix LABEL:app:pi0_libero.
@@
We evaluate real-world performance on a physical 6-DoF xArm6 arm controlled by OpenVLA-OFT. Figure 5 reports results under 75% token-pruning ratio. VLA-Pruner attains the highest relative accuracy, underscoring its practicality for on-robot deployment.
@@
We report inference efficiency metrics in Table 5, including inference latency ( or ) and FLOPs(T). At same pruning ratios, VLA-Pruner achieves comparable speedups to visual token–pruning methods. Its latency is 5 ms higher than FastV (due to distance computations in redundancy filtering), yet it achieves significantly better performance. At the 25% retention ratio, VLA-Pruner outperforms token pruning baselines at 50% retention, while being 16% faster and using 20% fewer FLOPs. VLA-Pruner is faster and more compute-efficient than VLA-Cache. VLA-Pruner requires storing historical attention, but it introduces a negligible increase in Max GPU memory (see Appendix LABEL:app:more_results:memory).
@@ 5.3 More Results
To validate our design, we evaluate three alternative token selection criteria: (1) prefill-only, relying solely on prefill attention; (2) action-only, utilizing only action decode attention; and (3) score-fusion, a naive weighted linear combination of both. As shown in Figure 6, prefill-only preserves high-level semantics but hinders fine-grained low-level control, whereas action-only improves short-horizon control but sacrifices semantic grounding and task planning, and thus neither maintains performance well.
Sensitivity Analysis
We analyze the sensitivity of VLA-Pruner to the decay window size in Figure 6, which demonstrates the robustness of our method. We further study varying decay rate and pruning layer in Appendix LABEL:app:more_results:ablations.
@@
We provide a qualitative case study of VLA-Pruner in Figure 7. VLA-Pruner preserves essential localized visual information for precise action execution, while maintaining broad semantic understanding.